Revise the data in your worksheet to reflect the results for the subsequent period as shown below:

Chapter 9: Applying Excel
Data
Revenue $16.50 q
Cost of ingredients $6.25 q
Wages and salaries $10,400
Utilities $800 + $0.20 q
Rent $2,200
Miscellaneous $600 + $0.80 q
Actual results:
Revenue $29,445
Cost of ingredients $11,425
Wages and salaries $10,605
Utilities $1,190
Rent $2,200
Miscellaneous $2,045
Planning budget activity 1,700 meals served
Actual activity 1,800 meals served

a. What is the activity variance for revenue? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance)).

b. What is the spending variance for the cost of ingredients? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance)).

c. What is spending variance for wages and salaries? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance)).

d. What is spending variance for total expenses? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance)).
Answer #1

a) Activity variance for revenue = (1700-1800)*16.50 = 1650 F

b) Spending variance for cost of ingredients = (1800*6.25-11425) = 175 U

c) Spending variance for wages and salaries = 10400-10605 =205 U

d) Spending variance for total expense = (1800*7.25+14000)-27465 = 415 U